A firm hires labor and capital to produce grapefruits. Currently the marginal product of the last unit of labor input is 40 and the marginal product of the last unit of capital input is 60. The market wage is $20, if the firm is using the optimal combination of inputs, then the price of capital is

Select one:

a. $30

b. $20

c. $100

d. $40
Answer #1

ANswer

An optimal level of output is produced at the per dollar marginal product of inputs should equal.

MPL/w=MPK/r

r=price of capital, w=wage, MPL=marginal product of capital , MPK=marginal product of labor

or

r=(MPK*w)/MPL

r=(60*20)/40

r=30

the price of capital is $30

option a